# Graded Betti numbers of the Jacobian algebra and total Tjurina numbers of plane curves

### v1

@@ Abstract.
In this paper we compute explicit closed formulas for the total Tjurina number $\tau(C)$ of a reduced projective plane curve $C$ in terms of the graded Betti numbers of the corresponding Jacobian algebra. This yields in particular a completely new view point on the classical upper bounds for the total Tjurina number $\tau(C)$ of a plane curve $C$ given by A. du Plessis and C. T. C. Wall.
@@ 2. Graded Betti numbers and partitions
with $c_{1}\leq...\leq c_{m-2}$ and $d_{1}\leq...\leq d_{m}$. It follows from [16, Lemma 1.1] that
@@
Using [16, Formula (13)], one obtains the relation
@@
for $j=1,\ldots,m-3$. It seems that the inequality in Remark 5.3 gives a new necessary condition on these Betti numbers.
@@ 3. Tjurina numbers for curves via graded Betti numbers of their Jacobian algebra
(i) Theorem 3.1 is a generalization of [2, Proposition 1.11] which treats the case $t=2$ and of [12, Theorem 3.1] which treats the case $t=3$.
@@ 4. Examples for curves of type $t(C)=4$
We recall first the following result due to du Plessis and Wall, see [13, Theorem 3.2] as well as [15] for an alternative approach.
@@
$\tau(C)-\tau(d,d_{1})_{min}=d_{1}^{2}+td_{2}-\frac{t^{2}+\sum_{j=1}^{n}\epsilon_{j}^{2}}{2}-\sum_{j=1}^{n}\epsilon_{j}d_{j+2}.$
However, the fact that this expression is positive is not obvious to us. It seems that the condition
$d_{1}^{2}+td_{2}-\frac{t^{2}+\sum_{j=1}^{n}\epsilon_{j}^{2}}{2}-\sum_{j=1}^{n}\epsilon_{j}d_{j+2}\geq 0$
coming from Theorem 5.1 is a necessary condition to be satisfied by the exponents $(d_{1},d_{2},\ldots,d_{m})$ and $(\epsilon_{1},\dots,\epsilon_{n})$, the corresponding ordered partition of $t$.
Recall that an $3$-syzygy curve $C$ is called nearly free (resp. minimal plus-one generated) if $d_{1}+d_{2}=d$ and $d_{2}=d_{3}$ (resp. $d_{1}+d_{2}=d$ and $d_{2}+1=d_{3}$, see [7, 10]. Using Corollary 5.2 (1) we can give new proofs for the following known results. Indeed, the claim (1) below occurs in [5, Theorem 1.3], and the claim (2) below occurs in [10, Theorem 1.5].
@@
Using Corollary 5.2 (2) we can give a completely new proof of [9, Theorem 3.1], that characterizes the maximal Tjurina curves in terms of their resolutions.
@@
from [11, Remark 4.3]. Then $C$ is a 4-syzygy curve with exponents $(6,6,6,7)$ and degree $d=11$. It follows that
@@
(ii) Using the final part of Remark 5.7 we have the following. The curves $C$ and $C^{\prime}$ in [11, Example 4.5] satisfy the conditions in Corollary 5.6 (2) for $d=7$, $d_{1}=4$ and $m=3$. The curves $C$ in [11, Example 4.6] satisfies the conditions in Corollary 5.6 (2) for $d=9$, $d_{1}=5$ and $m=3$. The curves $C$ in [11, Example 4.7] satisfies the conditions in Corollary 5.6 (2) for $d=11$, $d_{1}=6$ and $m=3$. To get examples of curves satisfying the conditions in Corollary 5.6 (2) with $m\geq 4$, one may consider the curves $C$, $^{\prime}, C′′′C^{\prime\prime\prime}$ and $C_{6}$ in [11, Example 5.1], the curves $C_{7}$ and $C^{\prime}$ in [11, Example 5.2], the curves $C$ and $C_{8}$ in [11, Example 5.3] and the curves $C^{\prime}$ and $C^{\prime\prime\prime}$ in [11, Example 5.4].

### v2

@@ Abstract.
In this paper we compute an explicit closed formula for the total Tjurina number $\tau(C)$ of a reduced projective plane curve $C$ in terms of the graded Betti numbers of the corresponding Jacobian algebra. This formula allows a completely new view point on the classical upper bounds for the total Tjurina number $\tau(C)$ of a plane curve $C$ given by A. du Plessis and C. T. C. Wall. This approach yields in particular a new necessary condition for a set of positive integers to be the graded Betti numbers of the Jacobian algebra of a reduced plane curve.
@@ 1. Introduction
This approach yields in particular a new necessary condition for a set of positive integers to be the graded Betti numbers of the Jacobian algebra of a reduced plane curve, see Remark 5.3.
@@ 2. Graded Betti numbers and partitions
with $c_{1}\leq...\leq c_{m-2}$ and $d_{1}\leq...\leq d_{m}$. It follows from [17, Lemma 1.1] that
@@
Using [17, Formula (13)], one obtains the relation
@@
for $j=1,\ldots,m-3$. We also know that
(2.6) $d_{1}\leq d_{2}\leq d_{3}\leq d-1,$
see for instance [9, Theorem 2.4] and
(2.7) $d_{m}\leq 2d-4,$
see for instance [6, inequality (2.6)]. In this note we prove the following inequality
(2.8) $E=d_{1}^{2}+td_{2}-\frac{t^{2}+\sum_{j=1}^{n}\epsilon_{j}^{2}}{2}-\sum_{j=1}^{n}\epsilon_{j}d_{j+2}\geq 0,$
which is a new necessary condition on these Betti numbers. In other words, (2.8) is not a consequence of the inequalities (2.5), (2.6) and (2.7), see for details Remark 5.3 below.
@@ 3. Tjurina numbers for curves via graded Betti numbers of their Jacobian algebra
(i) Theorem 3.1 is a generalization of [2, Proposition 1.11] which treats the case $t=2$ and of [13, Theorem 3.1] which treats the case $t=3$.
@@ 4. Examples for curves of type $t(C)=4$
We recall first the following result due to du Plessis and Wall, see [14, Theorem 3.2] as well as [16] for an alternative approach.
@@
$E=\tau(C)-\tau(d,d_{1})_{min}=d_{1}^{2}+td_{2}-\frac{t^{2}+\sum_{j=1}^{n}\epsilon_{j}^{2}}{2}-\sum_{j=1}^{n}\epsilon_{j}d_{j+2}.$
For $m=3$ the inequality $E\geq 0$ holds for obvious reasons, since one has $\epsilon_{1}=t$ and hence
$E=d_{1}(d_{1}-t)+t(d_{1}+d_{2}-t-d_{3})=d_{1}(d_{1}-t)+t(d-1-d_{3})\geq 0.$
Indeed, $d_{1}-t=d-1-d_{2}\geq 0$ and $d-1-d_{3}\geq 0$ by (2.6).
On the other hand, for $m\geq 4$, the inequality $E\geq 0$ coming from Theorem 5.1 is a new necessary condition to be satisfied by the exponents $(d_{1},d_{2},\ldots,d_{m})$ and $(\epsilon_{1},\dots,\epsilon_{n})$, the corresponding ordered partition of $t$.
To show that this inequality is not a consequence of the inequalities (2.5), (2.6) and (2.7), one may consider the data $d_{1}=3$, $d_{2}=d_{3}=8$, $d_{4}=15$, $\epsilon_{1}=\epsilon_{2}=1$ which verify the inequalities (2.5), (2.6) and (2.7), but not the above inequality. Indeed, for these data one gets E=1<0E=-1<0, which shows that there is no reduced curve $C$ with these Betti numbers.
Recall that an $3$-syzygy curve $C$ is called nearly free (resp. minimal plus-one generated) if $d_{1}+d_{2}=d$ and $d_{2}=d_{3}$ (resp. $d_{1}+d_{2}=d$ and $d_{2}+1=d_{3}$, see [8, 11]. Using Corollary 5.2 (1) we can give new proofs for the following known results. Indeed, the claim (1) below occurs in [5, Theorem 1.3], and the claim (2) below occurs in [11, Theorem 1.5].
@@
Using Corollary 5.2 (2) we can give a completely new proof of [10, Theorem 3.1], that characterizes the maximal Tjurina curves in terms of their resolutions.
@@
from [12, Remark 4.3]. Then $C$ is a 4-syzygy curve with exponents $(6,6,6,7)$ and degree $d=11$. It follows that
@@
(ii) Using the final part of Remark 5.7 we have the following. The curves $C$ and $C^{\prime}$ in [12, Example 4.5] satisfy the conditions in Corollary 5.6 (2) for $d=7$, $d_{1}=4$ and $m=3$. The curves $C$ in [12, Example 4.6] satisfies the conditions in Corollary 5.6 (2) for $d=9$, $d_{1}=5$ and $m=3$. The curves $C$ in [12, Example 4.7] satisfies the conditions in Corollary 5.6 (2) for $d=11$, $d_{1}=6$ and $m=3$. To get examples of curves satisfying the conditions in Corollary 5.6 (2) with $m\geq 4$, one may consider the curves $C$, $^{\prime}, C′′′C^{\prime\prime\prime}$ and $C_{6}$ in [12, Example 5.1], the curves $C_{7}$ and $C^{\prime}$ in [12, Example 5.2], the curves $C$ and $C_{8}$ in [12, Example 5.3] and the curves $C^{\prime}$ and $C^{\prime\prime\prime}$ in [12, Example 5.4].